Back to the NIKA2 Pool page

Daily Reports

October 30, (Tuesday)

Matthieu, Frédéric, Florian, Sean, Devika and Bilal are at the telescope.

Start of the pool... but it's snowing. We will wait until the weather clears up, which is expected around Thursday morning.

We are monitoring the KIDs behaviour in the dark for a while, waiting if there is any instabilities...

October 31, (Wednesday)

Same here, nothing could be done out of watching how stable the KIDs are when the window remain closed.

November 1, (Thursday)

We finally can observe with an incredible opacity of 0.150 and stable.

We had a few hiccups concerning the acquisition. The previous plan was to monitor any drift in calibration with the acquisition and see if it is correlated with anything tangible. But it turns out that last minute modification of the acquisition V2 rendered it useless, the pipeline being totally unable to process the raw data. Therefore, in a rush, with Juan online, we decided to switch immediately to v1. After a few adaptation, we could finally proceed with observation. NIKA2 IMBfits cannot be produced and therefore no monitoring with PIIC is possible this week. In these conditions it is impossible to have an estimate of the fluxes of the source we scan. This causes a problem for live follow-up of calibrators.

The following pattern is still used: we do science observation interrupted every two hours by scans on calibrators (MWC349, CRL2688 and NGC7027). During morning and afternoon, we followed the projects LP 192-16 and 050-18. The opacity remains under 0.3 but start to degrade after the sunset.

[Night shift : Frederic, Davika and Florian]

Opacity begins at 0.4 and keeps increasing during the night, unstable. We started the shift on 091-18. Around 21:45 UTC we began a set of calibration scans on Uranus : pointing, focus, and a beammap. We then produced 4 defocused beammaps for commissioning purposes, and repointed on Uranus. We checked our calibration using four secondary calibrators, and ran a skydip. We then went back to 091-18 for two hours with an opacity around .45, and follow with a new series of calibration scans, and tried to point on a source that unfortunately happened to be too faint to focus on. We then went back to 091-18. Around 04:30 UTC we had to retune the KIDS, and thus re-point the telescope.

November 2, (Friday)

bad weather - telescope stopped

[Night shift : Florian and Frédéric]

03:23 am : the wind speed came back to normal. We have restarted the NIKA2 acquisition. Tuning is not perfect but we can't do better for now. We start with a pointing on a bright source. It seems ok. We do a focus on the same source. It seems ok also.

3:55am : we have tried to retune the kids. Result is not that good...

4:00 : we do a beammap on Uranus

4:37 : skydip

5:00 : start project 091-18

November 3, (Saturday)

[Night shift : Florian and Frédéric]

10:00 pm starting the night shift. wind is ok. tau is high (0.4)

redo tuning

nikas-18 : pointing + Beammap on 3C84

10:54 : secondary calibrators

11:18 : beammap on uranus, then skydip

November 4, (Sunday)

[Night shift : Florian and Frédéric]

00:23 move to 091-18 project (Tony)

01:14 nikas-18 : restart DAQ, retune kids. focus on uranus

02:30 move to LPSZ : we try to do an ACT cluster (challenging) : LPSZ-ACT-CL-J0223.1-0056

03:30 back to nikas-18 to do a beammap (with "good" tau values)

04:57 : move to Deep Field. goodsn

[Night shift : Florian and Frédéric]

22:00 : nikas18. check pointing on uranus

22:24 : start LPSZ. PSZ2G155 cluster

We have done 3 scans and then the wind velocity increased (up to 80 km/h ! ) !!

Night shift might be already over...

November, 5 (Monday)

bad weather. Telescope stopped.

[Night shift : Florian]

Snow : could not observe until ~2AM UTC.

Retuned NIKA2 at 2AM UTC. Started observations at ~2:30 : pointing, focus and beammap on uranus.

The tau values seemed off (<0.1), which turned out to be due to a snow layer on the taumeter. After cleaning it, tau took more realistic values around 0.15.

02:45 : switched to 199-16 : pointing on 0336-019, 8 scans on PSZ2G193, and re-pointing on 0336-019.

03:30 : switched to nikas-18. Pointing & calib_1scan on 3C84, focus on 0316+413 (had to run three focuses because the first two did not work)

05:20 : switched to 192-16. Pointing on 1044+719 gives a FWHM(2mm) = 17.21 arcsec => OK according to the readme. Start scanning goodsnorth. After a full observation of Goodsnorth (i.e. 8 scans, the last one seems corrupted when looking at the output of the RTA ?), pointing on 1055+018, and then start scanning Cosmos (higher in the sky).

DailyReportsNika2Pool301018 (last edited 2018-11-06 05:52:10 by NikaBolometer)